# Anomalous Criticality of Absorbing State Transition toward Jamming

### v3

@@ Abstract
Abstract: The jamming transition is traditionally regarded as a geometric transition governed by static contact networks. Recently, dynamic phase transitions of athermal particles under periodic shear provide a new lens on this problem, leading to a conjecture that jamming transition corresponds to an absorbing-state transition within the Manna (conserved directed percolation) universality class. Here, by re-examining the biased random organization model, a minimal model for particles under periodic shearing that the conjecture is based on, we uncover several criticality anomalies at high density at odds with Manna universality class. In three-dimensional monodisperse systems, we find crystallization disrupts the absorbing transition, while in dense binary mixtures, a distinct transition from absorbing to active-glass states emerges, signifying a new universality class of dynamic phase transition. Closer to the jamming point, the quenched heterogeneity in the contact network smears the dynamic transition via Griffiths effects and drives the system toward heterogeneous directed percolation. We propose a field theory with fractional time dynamics that unifies these phenomena, establishing a theoretical framework linking jamming, disorder, and dynamic criticality.
@@ I Popular Summary
The jamming transition, when materials like sand or foam suddenly become rigid, is traditionally viewed as a geometric packing problem. Recent ideas suggest it’s also tied to a criticality of dynamic processes, like how crowded particles stop moving under repeated shaking or shearing. This work puts these ideas to the test and reveals a more complex and fascinating story: at intermediate high density, a new universality class emerges, marking a transition from absorbing to active-glass states, while near the jamming point, heterogeneity induces Griffiths effects smearing out the dynamic critical point. Our work reveals complicated interplay between static and dynamic phase transitions and sheds light on the dynamic criticality of high-dimensional disordered systems, like learning processes of artificial neural networks.
@@ II Introduction
The jamming transition, a transition from a mechanically soft state to a rigid state, widely exists in soft matter systems [trappe2001jamming, liu2010jamming, behringer2018physics, deng2024jamming]. The static packing of spheres provides a natural route to study the jamming transition, in which the random close packing (RCP) state is introduced to describe the most compact state that spheres can be arranged in a random way [bernal1960packing, scott1960packing, berryman1983random]. Although the RCP state can be protocol-dependent [torquato2000random, aste2005geometrical], it is still widely believed to be a signature of the jamming transition [danisch2010model, 2023Estimating, silbert2002geometry, skoge2006packing, 2004Jamming, donev2005unexpected, hexner2018two, shang2025jamming].
@@
Another path to tackle the jamming problem is from the dynamic side. It was first found that when dilute suspended colloidal particles are subjected to periodic shearing, a reversible to irreversible transition, or an absorbing phase transition, can occur upon increasing the shearing amplitude [pine2005chaos, corte2008random, reichhardt2023reversible]. As system density increases, the critical shearing amplitude becomes smaller, making the reversible to irreversible transition coincide with the jamming transition [schreck2013particle, milz2013connecting, reichhardt2023reversible, hima2014experimental, regev2013onset, regev2015reversibility, kawasaki2016macroscopic, nagasawa2019classification, pan2023review, deng2024jamming, ghosh2022coupled, das2020unified]. Nevertheless, directly studying the criticality of such a dynamic transition at high density is challenging, due to the divergent relaxation time and long-range correlation near the jamming point [regev2015reversibility, regev2013onset, reichhardt2023reversible, hima2014experimental]. Theoretically, the many-body dynamics of particles under periodic shearing can be well captured by a minimal model, i.e., the random organization (RO) model [corte2008random]. Recent works have demonstrated that the biased random organization (BRO) model, a more realistic version of the RO model, can generate the RCP state at the critical point of the absorbing transition when the displacement size (corresponding to shearing amplitude) vanishes $\varepsilon\rightarrow 0$ in dimensions $d=3$ to $5$ [wilken2021random, wilken2023dynamical, zhang2024absorbing]. The RO/BRO model in the dilute regime has been proven to belong to the Manna universality class, or conserved directed percolation (CDP) [menon2009universality, torquato2018hyperuniform, 2019Hydrodynamics, anand2025emergent, milz2013connecting]. In Ref. [wilken2021random, wilken2023dynamical, ness2020absorbing], it was suggested that the BRO model also resides in the Manna universality class in the close packing limit ($\varepsilon\rightarrow 0$). Nevertheless, a comprehensive and detailed study of the dynamic criticality of the BRO model in the high density regime is needed to fully resolve this issue.
@@
Figure 1: (a) Schematic of the BRO model: overlapping particles (red) become active and take pairwise repulsive displacements. (b) $f_{a}^{\infty}$ as a function of packing fraction $\varphi$ at $\varepsilon=2.0~\sigma$. (c) Evolution of $f_{b}(t)$ for systems at $\varepsilon=2.0~\sigma$ under various $\varphi$. The dashed line represents the power law $t^{-0.45}$. Here $N=65536$. (d) Critical packing fractions $\varphi_{c}$ for 2D binary (red line), 3D binary (orange dots), 3D monodisperse (blue circles) and 4D monodisperse (green open squares) systems at different $\varepsilon$. Asterisks mark the extrapolated $\varphi_{c}$ for $\varepsilon\rightarrow 0$, i.e., 0.840 (2D) and 0.638 (3D) and 0.455 (4D), which is consistent with the RCP.
@@
In this work, we show that the BRO model exhibits anomalous critical behaviors at high density, which deviates from Manna universality class. In 3D monodisperse sphere systems, we find that the absorbing transition is interrupted by crystallization. In binary mixture systems without crystallization, an unreported universality of dynamic phase transition, i.e., absorbing to active-glass transition, is identified, which is characterized by critical exponents different from the Manna universality. With further increasing the critical density by decreasing $\varepsilon$, dynamic heterogeneity and Griffiths effects emerge, which smooth the dynamic transition and broaden the critical point into a pseudo-critical regime. As $\varepsilon\rightarrow 0$, the dynamic transition becomes equivalent to heterogeneous directed percolation (DP) due to particle immobility set by the jamming. These results indicates that while the BRO model can generate RCP configurations, the criticality of the transition is fundamentally reshaped by the RCP state. This discourages a direct mapping of the jamming transition to an absorbing phase transitions. At last, we also propose a field theory with fractional time dynamics which can describe the transition from absorbing to active-glass transition and heterogeneous DP in the same framework, revealing a complicated interplay between dynamic criticality and spatial annealed/quenched disorder. Our findings can be a starting point to study the dynamic criticality of more complicate disordered systems, like the learning process of artificial neural networks [zhang2024absorbing, spigler2018jamming].
@@ III Model and simulations
In the BRO model, particles are first randomly distributed in the space of dimension $d$. In each discrete time step, if a particle is isolated without overlapping with other particles, it is considered inactive and remains stationary. On the contrary, if two particles overlap, they are considered as active and a pair of repulsive displacements with opposite direction are assigned to each other. The displacement pair has the same magnitude randomly chosen from uniform range of $0,\varepsilon/2]$, as shown in Fig. 1(a). If a particle overlaps multiple particles, pairwise displacements are superimposed [wilken2021random, wilken2023dynamical, zhang2024absorbing, galliano2023two]. We simulate $N$ spherical particles in a hypercubic box of length $L$ with periodic boundary conditions at packing fraction $\varphi$. Both the monodisperse particle system and equal-number binary mixture of small and large particles are considered. In the latter, the particle diameter ratio is $\sigma_{M}/\sigma_{m}=1:\sqrt{2}$ for 2D systems. For 3D systems, a larger particle diameter ratio $\sigma_{M}/\sigma_{m}=1:1.6$ is also considered. The small particle diameter in each system is chosen as the unit length $\sigma$.
@@
where $N_{a}(t)$ is the number of active particles at time step $t$. Here, $\tau=(\varphi-\varphi_{c})/\varphi_{c}$ is the distance to the critical point. Note that $_{a}$ here indicates that the average is done only for trials that remain active up to time $t$. Generally, below the critical point ($\tau<0$), the system falls into an absorbing phase with $f_{a}^{\infty}\equiv f_{a}(t\rightarrow\infty)=0$, while above the critical point ($\tau>0$), the system stays in an active state with $f_{a}^{\infty}>0$ (see Fig. 1(b)). At the critical point ($\tau=0$), $f_{a}(t)$ is expected to decay following a power law $t^{-\alpha}$. We can further define the overall activity of the system averaged over all trials
@@
where $P_{s}(\tau,t)$ is the survival probability of the active phase up to time $t$. At the critical point, $f_{b}(t)$ should also decay following a power law $t^{-\alpha}$. Below and above the critical point, $f_{b}(t)$ deviates from the power law as shown in Fig. 1(c). As shown later, $f_{b}(t)$ provides a good measure of the Griffiths effect. Details about the finite size scaling of $f_{a}(\tau,t)$, $f_{b}(\tau,t)$ and $P_{s}(\tau,t)$, the definition of critical exponents and other quantities can be found in the Methods section.
Figure 2: (a, d) Finite size scaling for survival activity $f_{a}(t)$ and survival probability $P_{s}(t)$ at $\varepsilon=0.8\sigma$, $\varphi_{c}=0.6324$, from which we obtain $\alpha=0.21,z=2.75$. (b, e) Finite size scaling for $f_{a}^{\infty}$ and $t^{*}$ at $\varepsilon=0.8\sigma$. (c, f) Critical exponents of $\beta$, $\nu_{\parallel}$, $\alpha$, $z$ as functions of $\varepsilon$, from which one can distinguish three regimes, I, II, III. Red and blue dotted lines represent the value of $\alpha,z,\nu_{\parallel},\beta$ in the Manna universality. (g, h) MSD(t) for systems with $\varepsilon=1.5\sigma$ (g) and $\varepsilon=0.5\sigma$ (h), where the open/solid symbols represent the absorbing/active state. (i) Diffusion coefficients $D$ (solid line) and $f_{a}^{\infty}$ (dotted line) as functions of $\varphi$ at various $\varepsilon$.
@@ IV Results
Based on the finite size scaling analysis, we can determine the critical packing fraction $\varphi_{c}$ of the system under different $\varepsilon$, which are summarized in Fig. 1(d). By extrapolating these data to $\varepsilon\rightarrow 0$, we obtain $\varphi_{c}=0.840$ in 2D, $\varphi_{c}=0.638$ in 3D, $\varphi_{c}=0.455$ in 4D for both monodisperse and binary systems. This is consistent with the reported RCP packing fractions $\varphi_{\rm RCP}$ [berryman1983random, silbert2002geometry, skoge2006packing, danisch2010model, 2023Estimating], and in line with previous studies [wilken2021random, wilken2023dynamical].
Figure 3: (a, e) Evolution of $f_{b}(t)$ at different $\varphi$, where Griffiths regime is painted in gray. (b, f) Evolution of $f_{b}(t)$ at different $N$. And $\varphi=0.8128$(upper) and $0.4920$ (lower). (c, g) Relationship between index $\kappa$ and $\varphi$. (d, h) Distributions of normalized activity duration $f_{ad}$ at $\varphi=0.7838,\varepsilon=0.2\sigma$(upper) and $0.4980,\varepsilon=1.5\sigma$, $\Delta\varepsilon=1.2\sigma$, $l=1\sigma$ (lower). (a, b) are for the original BRO model with $\varepsilon=0.1\sigma$. (e, f) are for modified BRO models with quenched-disorder at $\varepsilon=1.5\sigma$, $\Delta\varepsilon=1.2\sigma$ and $l=2\sigma$. For all systems except (b,f), $N=8192$.
@@
Nevertheless, for high density 3D monodispersed systems with $\varepsilon<0.8\sigma$, we find that the absorbing transition is disrupted by crystallization, thus does not exhibit standard critical phenomena (see Fig. S1). Although binary systems can avoid crystallization, their critical behaviors do not match the Manna universality. In Fig. 2(a, b, d, e), we show the finite-size scaling of $f_{a}(t)$, $P_{s}(t)$, $f_{a}^{\infty}(\tau)$ and $t^{*}(\tau)$ for 2D binary systems with $\varepsilon=0.8\sigma$ at $\varphi_{c}=0.6324$. This finite-size scaling is different from the typical BRO model with a large step size ($\varepsilon=1.5\sigma$) (see Fig. S2). The obtained critical exponents $\beta$, $\nu_{\parallel}$, $\alpha$, $z$ as functions of $\varepsilon$ are shown in Fig. 2(c, f), where for $\varepsilon>1.0\sigma$ (Regime I), these critical exponents are consistent with the Manna universality. When $0.3\sigma<\varepsilon<1.0\sigma$ (Regime II), we find non-monotonic behaviors in $\nu_{\perp}$, $\alpha$, and $z$ inconsistent with the Manna universality. Further decreasing $\varepsilon$ below $0.3\sigma$ (Regime III), the failure of standard finite size scaling and the uncertainty of the critical point are observed. It should be noted that the exponent $\beta$ does not change significantly in Regimes I and II. The measured $z$, $z^{\prime}$, and $z^{*}$ are consistent, and $\nu_{\perp}=z/\nu_{\parallel}$ also changes with $\varepsilon$ (see Fig. S3). In Fig. S3-6, we also present the exponents measured in the 3D system under two different particle diameter ratios, as well as the monodisperse 4D system. We find that the general behavior of the critical exponent and finite size scaling remain unchanged. This indicates that the anomalous criticality of the BRO model is robust at high density, insensitive to the dimensionality and particle diameter ratio.
@@
We first focus on the Regime I where the step size $\varepsilon$ is larger than the particle size. In this regime, the motion of active particles is not constrained by neighbouring particles. Thus, the mean-square-displacement (MSD) of active particles shows diffusion scaling as shown in Fig. 2(g). Here, the solid and open symbols represent systems above and below the critical point, respectively. The diffusion scaling in sub-critical systems indicates that the absorbing transition is essentially a non-local searching process for a non-overlapping configuration, consistent with Manna (CDP) universality [Manna_1991, wiese2024hyperuniformity].
As $\varepsilon$ decreases, $\varphi_{c}$ increases. Due to the repulsive displacement in the BRO model, the particles are caged by their neighbours. This effect is shown in Fig. 2(h) for systems with $\varepsilon=0.5\sigma$, where one can find an early-time plateau or sub-diffusion regime in MSD for systems near the critical point, a characteristic of the glass-forming liquids [debenedetti2001supercooled, berthier2011theoretical, 2025Structural, tanaka2010critical]. In Fig. 2(i), we plot the measured diffusion coefficient $D$ of active particles and $f_{a}^{\infty}$ as a function of $\varphi$ under different $\varepsilon$. We can clearly distinguish active states with $D$ as low as $10^{-6}\sigma^{2}/\tau_{0}$ for $\varepsilon\lesssim\sigma$. The decoupling between $f_{a}^{\infty}$ and $D$ suggests that in this regime (Regime II), the absorbing phase transition is essentially an absorbing to active-glass transition, where particles can only find a non-overlapping absorbing configuration through local arrangements instead of non-local searching.
The fundamental difference between these two scenarios can be distinguished by field theory. For Manna (CDP) universality, it can be described by a non-conserved field $\rho({\bf{r}},t)$ (active particle density field) coupling with a non-diffusive conserved field $\psi({\bf{r}},t)$ (total density field) [rossi2000universality, pastor2000field], namely,
@@
with $\theta=1$ indicating the evolution of $\psi(t)$ field is a result of diffusion of $\rho(t)$ field. Note that $a$, $b$, $w$ are fixed parameters determining the critical point [pastor2000field]. The last term in Eq.(3) is the multiplicative noise term accounting for absorbing phase transition. In the active-glass state, however, the dynamic of $\psi(t)$ field in Eq.(4) is sub-diffusive due to cage effect, corresponding to fractional time derivative with $0<\theta<1$ [metzler2000random, metzler1999anomalous], while the activity field is still diffusive as activity can spread by particle contact. Obviously, the criticality with fractional time dynamics distinguishes itself from Manna universality. The observed continuous variation of critical exponents in this regime might be due to the variation of $\theta$. Nevertheless, numerical solution and renormalization-group analysis of this fractional time equation with multiplicative noise currently remains challenging [metzler2000random, metzler1999anomalous]. It should be noted that a field description of glass state with a fractional time derivative is a highly simplified treatment. Further theoretical explorations are needed to confirm whether this simplification preserves all relevant factors near the critical point.
Griffiths effect in pre-jamming state
When $\varepsilon$ is further reduced below $0.3\sigma$ (Regime III), the critical density of the system would approach that of RCP state. In this regime, caging-breaking events become unlikely and the topology of particle network remains unchanged in the activation process. We refer to this state as the pre-jamming state. Different from the conventional absorbing phase transition, we find $f_{b}(t)$ in this state relaxes according to a continuous power law
@@
with the $\varphi$-dependent critical exponent

### v4

@@ Abstract
Abstract: Jamming transition is traditionally regarded as a geometric transition governed by static contact networks. Recently, dynamic phase transitions of athermal particles under periodic shearing provide a new lens on this problem, leading to a conjecture that jamming transition corresponds to an absorbing-state transition within the Manna (conserved directed percolation) universality class. Here, by re-examining biased random organization models, minimal models for particles under periodic shearing that the conjecture is based on, we uncover several criticality anomalies at high density at odds with the Manna universality class. In three-dimensional monodisperse systems, we find crystallization disrupts the absorbing transition, while in dense binary mixtures, a distinct transition from absorbing to active-glass state emerges, signifying a new dynamic universality class. Close to the jamming point, the quenched heterogeneity in the contact network of binary systems smears the dynamic criticality via Griffiths effects and drives the system toward heterogeneous directed percolation. For close-packed crystal structures, Griffiths effect is absent. However, the dynamic criticality still deviates from the Manna model. These phenomena are explained by a field theory with fractional time dynamics that links jamming, disorder and dynamic criticality.
@@ I Popular Summary
Jamming transition, when materials like sand or foam suddenly become rigid, is traditionally viewed as a geometric packing problem. Recent ideas suggest it’s also tied to a criticality of dynamic processes, like how crowded particles stop moving under repeated shaking or shearing. This work puts these ideas to the test and reveals a more complex and fascinating story: at intermediate high density, a new universality class of dynamic transition from absorbing to active-glass state emerges, while near the jamming point, heterogeneity induces Griffiths effects smearing out the dynamic critical point. Our work reveals complicated interplay between static and dynamic phase transitions and sheds light on the dynamic criticality of high-dimensional disordered systems, like learning processes of artificial neural networks.
@@ II Introduction
The jamming transition, a transition from a mechanically soft state to a rigid state, widely exists in soft matter systems [75, 37, 4, 16]. The static packing of spheres provides a natural route to study the jamming transition, in which the random close packing (RCP) state is introduced to describe the most compact state that spheres can be arranged in a random way [5, 64, 6]. Although the RCP state can be protocol-dependent [72, 3], it is still widely regarded as a signature of the jamming transition [13, 2, 66, 67, 18, 17, 28, 65].
@@
Another path to tackle the jamming problem is from the dynamic side. It was first found that when dilute suspended colloidal particles are subjected to periodic shearing, a reversible to irreversible transition, or an absorbing state transition, can occur upon increasing the shearing amplitude [55, 12, 59]. As system density increases, the critical shearing amplitude becomes smaller, making the reversible to irreversible transition coincide with the jamming transition [63, 45, 59, 30, 57, 58, 33, 48, 52, 16, 20, 14]. Nevertheless, directly studying the criticality of such a dynamic transition at high density is challenging, due to the divergent relaxation time and long-range correlation near the jamming point [58, 57, 59, 30]. Theoretically, the many-body dynamics of particles under periodic shearing can be well captured by a minimal model, i.e., the random organization (RO) model [12, 71, 26]. Recent works have demonstrated that the biased random organization (BRO) model, a more realistic version of the RO model in which the displacements of particles are not fully random but repulsive, can generate the RCP state at the critical point of the absorbing transition when the displacement size (corresponding to shearing amplitude) vanishes $(\varepsilon\rightarrow 0)$ in dimensions $d=1$ to $5$ [84, 85, 87]. The RO/BRO model in the dilute regime has been proven to belong to the Manna universality class, or conserved directed percolation (CDP) [43, 74, 36, 1, 45]. In Ref. [84, 85, 49], it was suggested that the BRO model also resides in the Manna universality class in the close-packed limit ($\varepsilon\rightarrow 0$). Nevertheless, for Manna model and directed percolation under $d<4$, the Harris criterion ($d\nu_{\perp}>2$)[23, 76, 78, 77] is violated. This implies that disorder may have significant influence on the dynamic criticality. Therefore, a comprehensive and detailed study of the dynamic criticality of the BRO model in the high density regime is needed to fully resolve this issue.
@@
Figure 1: (a) Schematic of conserved BRO model: overlapping particles (red) become active and take pairwise repulsive displacements. (b) Order parameter $f_{a}^{\infty}$ as a function of packing fraction $\varphi$ at $\varepsilon=2.0\sigma$. (c) Evolution of $f_{b}(t)$ for systems at $\varepsilon=2.0\sigma$ under various $\varphi$. The dashed line represents the power law $t^{-0.42}$. Here $N=65536$. (d) Critical packing fractions $\varphi_{c}$ for 2D binary (red line), 3D binary (orange dots), 3D monodisperse (blue circles) and 4D monodisperse (green open squares) systems with conserved BRO dynamics at different $\varepsilon$. Here $\overline{\sigma}$ represents the average diameter of particles. Asterisks mark the extrapolated $\varphi_{c}$ for $\varepsilon\rightarrow 0$, i.e., 0.838 (2D), 0.645 (3D binary), 0.640 (3D monodisperse) and 0.453 (4D), which are consistent with the RCP states [6, 66, 67, 13, 2].
@@
In this work, we study two types of BRO model with or without center of mass conservation (CMC) [27, 84, 85] and find both of them exhibit anomalous critical behaviors at high density, which deviate from the Manna universality class. In 3D monodisperse sphere systems, we find that the absorbing transition is interrupted by crystallization. In binary mixture systems without crystallization, an unreported universality class of dynamic phase transition, i.e., absorbing to active-glass transition, is identified, which is characterized by critical exponents different from the Manna universality class. With further increase of the critical density by decreasing $\varepsilon$, dynamic heterogeneity and Griffiths effects emerge, which smooth the dynamic transition and broaden the critical point into a pseudo-critical regime. As $\varepsilon\rightarrow 0$, the dynamic transition becomes equivalent to heterogeneous directed percolation (DP) due to particle immobility set by the jamming. These results indicate that while the BRO models can generate RCP configurations, the criticality of the transition is fundamentally reshaped by the RCP state. This discourages a direct mapping of the jamming transition to an absorbing phase transition for disordered systems under $d\leq 4$. For closed-packed crystal structures, the Griffiths effect is absent. The dynamic criticality is thus restored, but still deviates from the Manna universality class. We propose a field theory with fractional time dynamics which can unify the above phenomena in the same framework, revealing a complicated interplay between dynamic criticality and spatial annealed/quenched disorder. Our findings can be a starting point to study the dynamic criticality of more complicated disordered systems, like the learning process of artificial neural networks [87, 68, 1].
@@ III Model and simulations
In BRO models, particles are first randomly distributed in the space of dimension $d$. In each discrete time step, if a particle is isolated without overlapping with other particles, it is considered inactive and remains stationary. On the contrary, if two particles overlap, they are considered as active and a pair of repulsive central displacements with opposite directions are assigned to each other, as shown in Fig. 1(a) [27, 84, 85, 87, 19]. In this work, we explicitly consider two types of BRO dynamics, classified by whether they conserve the center of mass of the system [27]. In the conserved dynamics, the displacements of a pair of overlapped particles have the same magnitude randomly chosen from a uniform range of $0,\varepsilon/2]$. If a particle overlaps with multiple particles, these displacements (vectors) are superimposed [27, 19]. In the non-conserved dynamics, the displacement magnitude of each overlapped particle is randomly chosen from $0,\varepsilon/2]$, and the direction of displacement of the particle is from the mass center of overlapped clusters to the center of this particle [84, 85]. Since the two types of BRO model exhibit similar results, in the main text, we mainly show the results of conserved BRO model, while leaving the data of non-conserved one in Supplementary Information (SI).
We simulate $N$ spherical particles in a hypercubic box of length $L$ with periodic boundary conditions at packing fraction $\varphi$. Both monodisperse particle systems and equal-number binary mixtures of small and large particles are considered. In the latter, the particle diameter ratio is $\sigma_{M}/\sigma_{m}=1:\sqrt{2}$. The small particle diameter in each system is chosen as the unit length $\sigma$ and $\tau_{0}$ is the unit time.
@@
where $N_{a}(t)$ is the number of active particles at time step $t$. Here, $\tau=(\varphi-\varphi_{c})/\varphi_{c}$ is the distance to the critical point. Note that $_{a}$ here indicates that the average is only on trials that remain active up to time $t$. Generally, below the critical point ($\tau<0$), the system falls into an absorbing phase with $f_{a}^{\infty}\equiv f_{a}(t\rightarrow\infty)=0$, while above the critical point ($\tau>0$), the system stays in an active state with $f_{a}^{\infty}>0$ (see Fig. 1(b)). At the critical point ($\tau=0$), $f_{a}(t)$ is expected to decay following a power law $t^{-\alpha}$.
We can further define the overall activity of the system averaged over all trials
@@
where $P_{s}(\tau,t)$ is the survival probability of the active phase up to time $t$, which can be used to define the characteristic life time $t^{*}$ of the trials. At the critical point, $f_{b}(t)$ should also decay following a power law $t^{-\alpha}$. Below and above the critical point, $f_{b}(t)$ deviates from the power law (see Fig. 1(c)). As shown later, $f_{b}(t)$ provides a good measure of the Griffiths effects. Details about the finite-size scaling of $f_{a}(\tau,t)$, $f_{b}(\tau,t)$, $P_{s}(\tau,t)$ and $t^{*}(\tau)$, the definition of critical exponents and other quantities can be found in the Methods Section.
@@ IV Results
Based on the finite-size scaling analysis, we can determine the critical packing fraction $\varphi_{c}$ of systems under different $\varepsilon$, which are summarized in Fig. 1(d). By extrapolating these data to $\varepsilon\rightarrow 0$, we obtain $\varphi_{c}=0.838$ in 2D, $\varphi_{c}=0.645$ in 3D binary, $\varphi_{c}=0.640$ in 3D monodisperse, $\varphi_{c}=0.453$ in 4D for both monodisperse and binary systems. Similar data for non-conserved BRO model can be found in Fig. S1. These values are consistent with the reported RCP packing fractions $\varphi_{\rm RCP}$ [6, 66, 67, 13, 2], and in line with previous studies [84, 85].
Figure 2: (a) The evolution of the activity $f_{b}(t)$ at different packing fractions at $\varepsilon=0.5\sigma$. (b) The particle fraction in crystal phase $\rho_{cr}$ under different $\varepsilon$ and packing fractions at $t=10^{6}\tau_{0}$. Inset: The crystallization in a 3D monodisperse system under conserved BRO dynamic with $\varepsilon=0.5\sigma$ at packing fraction $\phi=0.480$ and $t=10^{6}\tau_{0}$. Different colors represent different crystal grains.
Nevertheless, for 3D monodisperse conserved BRO model under $0.2\sigma<\varepsilon<0.7\sigma$, we find that the dynamic transition does not exhibit standard critical phenomena (Fig. 2(a)). Detailed structural analysis based on classical crystal order parameter (see SI) reveals that the dynamic transition is disrupted by crystallization (Fig. 2(b)). For non-conserved BRO model, crystallization is also observed when $0.2\sigma<\varepsilon<0.5\sigma$ (See Fig. S2). This forces us to focus on binary system, where crystallization is avoided. However, we find that the critical behaviors of binary systems still mismatch the Manna universality class.
Figure 3: (a, c) Finite size scaling for survival probability $P_{s}(t)$ and survival activity $f_{a}(t)$ at $\varepsilon=0.8\sigma$, $\varphi_{c}=0.6496$. (b, d) Finite size scaling for $t^{*}$ and $f_{a}^{\infty}$ at $\varepsilon=0.8\sigma$. (e, f, g, h) Critical exponents of $z^{*}$, $\nu_{\parallel}$, $\alpha$, $\beta$ and $\nu_{\perp}$ as functions of $\varepsilon$, from which one can distinguish three regimes (from right to left): regime I (light blue), regime II (light orange), regime III (light red). The dotted lines represent the values of $z^{*},\nu_{\parallel},\alpha,\beta$ and $\nu_{\perp}$ in the Manna universality class [39, 25].
@@
In Fig. 3(a-d), we show the finite-size scaling of $f_{a}(t)$, $P_{s}(t)$, $f_{a}^{\infty}(\tau)$ and $t^{*}(\tau)$ for 2D binary systems with $\varepsilon=0.8\sigma$ at $\varphi_{c}=0.6496$. The obtained critical exponents $\beta$, $\nu_{\parallel}$, $\alpha$, $z^{*}$ and $\nu_{\perp}$ as functions of $\varepsilon$ are shown in Fig. 3(e-h). For $\varepsilon>1.0\sigma$ (Regime I, cyan), these critical exponents are consistent with the Manna universality class. When $0.3\sigma<\varepsilon<1.0\sigma$ (Regime II, orange), we find non-monotonic behaviors of $\nu_{\perp}$, $\alpha$, and $z$, inconsistent with the Manna universality class. Further decreasing $\varepsilon$ below $0.3\sigma$ (Regime III, red), the failure of standard finite size scaling and the uncertainty of the critical point are observed (In Fig. S3-S5). In Fig. S6-10, we also present the exponents measured in three and four dimensional systems for both conserved and non-conserved BRO systems. We find that the general behavior of the critical exponents and finite size scaling remains qualitatively the same. This indicates that the anomalous criticality of the BRO model is robust at high density, insensitive to the dimensionality, particle ratio and CMC.
@@
In Regime I, the step size $\varepsilon$ is larger than the particle size. Thus, the motion of active particles is not constrained by neighbouring particles. Thus, the mean-square-displacement (MSD) of active particles shows diffusion scaling as shown in Fig. 4(a). Here, the solid and open symbols represent systems above and below the critical point, respectively. The diffusion scaling in sub-critical systems indicates that the absorbing transition is essentially a non-local searching process for a non-overlapping configuration, a characteristic of the Manna universality class [41, 83].
As $\varepsilon$ decreases, $\varphi_{c}$ increases. Due to the repulsive displacement in the BRO model, the particles are caged by their neighbours. This effect is shown in Fig. 4(c) for systems with $\varepsilon=0.5\sigma$, where one can find an early-time plateau or sub-diffusion regime in MSD for systems above the critical point, a characteristic of the glass-forming liquids [15, 7, 70, 69]. In Fig. 4(b), we plot the measured diffusion coefficient $D$ of active particles and $f_{a}^{\infty}$ as a function of $\varphi$ under different $\varepsilon$. We can clearly distinguish active states with $D$ as low as $10^{-6}\sigma^{2}/\tau_{0}$ for $\varepsilon<\sigma$. The decoupling between $f_{a}^{\infty}$ and $D$ suggests that in Regime II, the absorbing phase transition is essentially an absorbing to active-glass transition, where particles can only find a non-overlapping absorbing configuration through local arrangements instead of non-local searching. Interestingly, when $\varphi$ further increases, the system with $\varepsilon<1.0\sigma$ becomes diffusive again, which corresponds to the yielding transition of granular systems under oscillatory shearing [58, 59, 30, 33, 48]. In Fig. 4(d), we summarize the dynamic behaviors of the system in a phase diagram, where the active-glass phase can be clearly identified in the cyan regime. The non-conserved BRO model also has a similar phase diagram (see Fig. S11).
Figure 4: (a, c) MSD(t) for systems under different $\varepsilon$, where the open/solid symbols represent the absorbing/active state, respectively. (b) Diffusion coefficients $D$ (solid line) and $f_{a}^{\infty}$ (dotted line) as functions of $\varphi$ at various $\varepsilon$. (d) Dynamic phase diagram of the system in the dimensions of packing fraction $\varphi$ and $\varepsilon$. The yellow line is the absorbing-active transition line. The blue line indicates the ’yielding transition’ line (cyan regime).
Figure 5: Griffiths effects for original BRO model with $\varepsilon=0.1\sigma$ (upper row) and modified BRO models with quenched-disorder ($\varepsilon=1.5\sigma$, $\Delta\varepsilon=1.2\sigma$) (lower row). (a, e) Evolution of $f_{b}(t)$ at different $\varphi$, where Griffiths regimes are painted in gray. Here, $N=8192$. (b, f) Evolution of $f_{b}(t)$ at different $N$. Here, $\varphi=0.8152$ (upper) and $0.5160$ (lower). (c, g) Distributions of normalized activity duration $f_{ad}$ for systems studied in (b, f) with $N=8192$. (d, h) Measured $\kappa$ index (Eq.(5)) as a function of $\varphi$ under different $\varepsilon$ and $l$. In (d), $\varphi^{+}_{c}$ = $0.7916$, $0.8170$, $0.8297$ for $\varepsilon=0.2$, $0.1$, $0.05\sigma$, respectively. In (h), $\varphi^{+}_{c}$ = $0.526$, $0.527$, $0.531$ for $l=1,2,4\sigma$, respectively.
The fundamental difference between absorbing to active-glass transition and the Manna universality class can be distinguished by field theory. It is known that the Manna universality class is described by a non-conserved field $\rho({\bf{r}},t)$ (active particle density field) coupling with a non-diffusive conserved field $\psi({\bf{r}},t)$ (total density field) [62, 53], namely,
@@
with $\theta=1$. Note that $a$, $b$, $w$ are fixed parameters determining the critical point [53]. The last term in Eq.(3) is the multiplicative noise term accounting for absorbing state transition. In the active-glass state, however, the dynamics of the $\psi(t)$ field in Eq.(4) is sub-diffusive due to the caging effect. This effect causes long-range temporal correlation, which can be phenomenologically described by fractional time derivative with $0<\theta<1$ [44, 9, 29, 42, 82, 61], while the activity field is still diffusive as activity can spread by particle contact. Note that in principle one can also construct more sophisticated theories based on the mode-coupling theory [60, 22] to describe the glassy evolution of $\psi(t)$. Nevertheless, the fractional time description gives a more intuitive picture of how the caging effect changes the coupling between $\psi(t)$ and $\rho(t)$ fields, thus modifies the criticality of the Manna universality class. Further theoretical explorations are needed to confirm whether this simplification preserves all relevant factors near the critical point.
Griffiths effects in pre-jamming state
When $\varepsilon$ is further reduced below $0.3\sigma$ (Regime III), the critical density of the system would approach that of RCP state. In this regime, caging-breaking events become unlikely and the topology of particle network remains unchanged in the activation process. We refer to this state as the pre-jamming state. Different from the conventional absorbing state transition, we find $f_{b}(t)$ in this state relaxes according to a continuous power law
@@
with the $\varphi$-dependent index [76]